The area of a rectangle is given by the following polynomial:
44t - 2t^2
If the length of one side is 2t, what is the polynomial that expresses the length of the other side?

Found 2 solutions by dabanfield, scott8148:
Answer by dabanfield(803)

The area of a rectangle is given by the following polynomial:
44t - 2t^2
If the length of one side is 2t, what is the polynomial that expresses the length of the other side?
Area = length * width so:
44t - 2t^2 = L * 2t
We have then:
L = (44t - 2t^2)/2t
L = [2t*(22-t)]/2t
L = 22 - t

Answer by scott8148(6628)

(44t - 2t^2) / 2t = 22 - t